# Shock waves in nonlinear transmission lines

### v4

@@ Abstract
In the first half of the paper we consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate the ”sound” wave coefficient of reflection from (coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the speeds of the ”sound” waves relative to the shock and the wave impedances. In the second half of the paper we explicitly include into consideration the dissipation in the system, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to justify the conditions on the shocks, postulated in the first half of the paper. This also allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
@@ V Lossy transmission line
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const},$ (38a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const},$ (38b)
To study the shock waves we should impose the boundary conditions
@@
Equating the r.h.s. of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a). Doing the same procedure with (38b) we recover the jump conditions (10b). Because the quantities ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$), satisfy the constraint (21), Eq. (39) specifies 3 boundary conditions, which thus define also the shock speed $U$. Continuing the analysis of the boundary conditions, we notice that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (42)
@@
$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{1}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{1}}\end{array}\right|<0.$ (46)
@@
$U_{21}^{2}>u_{1}^{2}.$ (47)
Similarly for the point $(Q_{2},\Phi_{2})$ we obtain
$u_{2}^{2}>U_{21}^{2}.$ (48)
@@
Note that the kinks, which we considered in our previous publications in addition to shocks but do not consider here, correspond to both fixed points being semistable. Thus the kinks turn out to be supersonic both with respect to the phase before the front and the phase behind the front.
V.1 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section.. We can exclude the current from (38) and after a bit of algebra obtain
@@ V.1 Half-nonlinear transmission line
To (49) we should add the boundary conditions (39a). (The third arbitrary constant appears when we connect $\Phi$ with the found $Q$.)
Equation (49) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (50). The boundary conditions (39a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$. We may present these maximum/minimum conditions as
@@
and opposite inequality for $V=V_{1}$. Recalling (35) we realise that we have reproduced in this particular case Eqs. (47), (48).
V.1.1 Weak shocks
Before going further it is convenient to rewrite (50) as
@@ V.1.1 Weak shocks
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
@@
where $\beta$ and $C_{s}$ are assumed to be constant. We also may take $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. So finally Eq. (49) becomes
@@
$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}},$ (55a)
$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}}.$ (55b)
One can check up that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(Q_{1}-Q_{2})}{6},$ (56)
there exists the elementary solution of (54)
@@
$\chi=\frac{\tau_{R}}{5T^{2}}.$ (58)
Approximation (53) can be inadequate, if $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (53) to
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}}.$ (59)
In this case (49) becomes
where $Q_{3}=\beta/\gamma+Q_{1}+Q_{2}$. One can check up that when the parameters of the problem satisfy the condition
there exists the elementary solution of (60)
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (62)
where
$\psi=\frac{\tau_{R}\left(Q_{1}-Q_{2}\right)}{T^{2}\left(Q_{1}+Q_{2}+2Q_{3}\right)}.$ (63)
In the Appendix A we’ll show how one arrives to the solutions (57) and (62).
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems and Workshop on Localization and Ergodicity, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A The ODE which doesn’t contain explicitly the independent variable
Equation (49) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $Q$ as the new independent variable and
${\cal I}=\frac{dQ}{d\tau}$ (64)
as the new dependent variable. In the new variables (49) takes the form of Abel equation of the second kind polyanin
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=F(Q).$ (65)
As it follows from (39a), the boundary conditions for (65) are
$\lim_{Q\to Q_{1},Q_{2}}{\cal I}(Q)=0.$ (66)
Since (65) is the first order differential equation, the problem is overdetermined. However, for the problems considered explicitly in Section V.1.1, not only will we be able to find the solutions of (65) satisfying the boundary conditions (66) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
A.1 Integration in terms of elementary functions
For $Q(V)$ given by (59), Eq. (65) can be presented as
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}).$ (67)
The boundary conditions (66) and the r.h.s. of (67) prompt the idea kogan4 ; kogan5 to look for ${\cal I}(Q)$ as the polynomial function proportional to $\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$. Because the r.h.s. of (67) is the third order polynomial of $V$, the multiplying factor should be a constant
${\cal I}(Q)=\frac{\psi}{Q_{1}-Q_{2}}\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$ (68)
(the normalization was chosen for later convenience).
Substituting (68) into (67) we obtain
$\frac{T^{2}\psi^{2}}{\left(Q_{1}-Q_{2}\right)^{2}}\left(2Q-Q_{1}-Q_{2}\right)+\frac{\tau_{R}\psi}{Q_{1}-Q_{2}}=\gamma\left(Q+Q_{3}\right).$
Hence the condition on the parameters of the problem, for (68) to be the solution of (67) is given by (61), and the constant multiplier in (68) should is given by (63). Substituting (68) into Eq. (64) and integrating we obtain the solution (62).
For $Q(V)$ given by (53) we obtain, instead of (67),
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (70)
If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial equal to zero at $Q=Q_{1,2}$. So let us split in two one multiplier in the solution (68)
$\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
This prompts us the idea to look up for the solution of (70) in the form kogan4 ; kogan5
${\cal I}(Q)=2\chi\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left(Q-Q_{2}\right).$ (72)
Substituting the ansatz into (70) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (56) and $\chi$ is given by (58). Substituting thus obtained ${\cal I}(Q)$ into (64) and integrating we obtain (57)
A.2 Solitary wave solution of a generalized KdV equation
To additionally illustrate the method of integration used in this Appendix, let us recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
@@ A.2 Solitary wave solution of a generalized KdV equation
Appendix B Shock profile for the fully nonlinear transmission line
For the fully nonlinear transmission line it is convenient to present the problem in the form slightly different from that used everywhere above. We’ll consider the shock speed $U$ and the conditions before the shock as given. Thus the boundary conditions are
$\lim_{\tau\to-\infty}Q(\tau)=Q_{1};\hskip 14.22636pt\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1}$ (82)
and we can concretize (38) to
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=U(Q-Q_{1})-I+I(\Phi_{1}),$ (83a)
$R_{C}\frac{dQ}{d\tau}$ $=U(\Phi-\Phi_{1})-V+V(Q_{1}).$ (83b)
Equation (83) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (83) either $R_{C}=0$ or $R_{L}=\infty$. In the first case we equate the r.h.s. of (83b) to zero and obtain
$V=U(\Phi-\Phi_{1})+V(Q_{1}).$ (84)
$I=U(Q-Q_{1})+(\Phi_{1}).$ (85)
If both $R_{C}>0$ and $R_{L}<\infty$, from (83) we may obtain the equation
$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U(\Phi-\Phi_{1})-V(Q)+V(Q_{1})}{U(Q-Q_{1})-I(\Phi)+I(\Phi_{1})}.$ (86)
If we manage to solve (86), we can substitute the solution into one of the equations (83) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (54) and (60) describing weak shocks in half-nonlinear transmission line (i.e. Eqs. (56) and (61)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (86) we postpone until later.

### v5

@@ Abstract
In the first half of the paper we consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate the ”sound” wave coefficient of reflection from (coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the speeds of the ”sound” waves relative to the shock and the wave impedances. In the second half of the paper we explicitly include into consideration the dissipation in the system, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to justify the conditions on the shocks, postulated in the first half of the paper. This also allows us to describe the shocks as physical objects of finite width and study their profiles, same as the profiles of the waves closely connected with the shocks - the kinks. The profiles of the latter, and in some particular cases the profiles of the former, were obtained in terms of elementary functions.
@@ V Lossy transmission line
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const}_{1},$ (38a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const}_{2},$ (38b)
Equation (38) contains 3 arbitrary constants - $U,\text{const}_{1},\text{const}_{2}$ hence to define the solution unequivocally we need 5 boundary conditions. Because we are ready to consider the solution up to translation in time, 4 of those would be enough. Since we are considering localized travelling waves, it would be convenient to take as those the quantities $Q_{1},Q_{2},\Phi_{1},\Phi_{2}$, defined by the equations
@@
Equating the r.h.s. of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a). Doing the same with (38b) we recover the jump conditions (10b).
The boundary conditions show that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point (see, however, Section VIII.1). (Other options would have demanded fine tuning of the parameters.) This means
$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{2},\Phi_{2}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{2},\Phi_{2}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{2},\Phi_{2}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{2},\Phi_{2}}\end{array}\right|>0,$ (42)
@@
$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{2}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{2}}\end{array}\right|>0.$ (46)
@@
$U_{21}^{2}<u_{2}^{2}.$ (47)
Similarly for the point $(Q_{1},\Phi_{1})$ we obtain
$u_{1}^{2}<U_{21}^{2}.$ (48)
@@
VI Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section. We can exclude the current from (38) and after a bit of algebra obtain closed equation for $Q(\tau)$
@@ VI Half-nonlinear transmission line
The solution of (49) should satisfy the boundary conditions (39a). (The third constant appears when we connect $\Phi$ with the found $Q$.)
Equation (49) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well $\Pi(Q)$ defined by the equation
$-\frac{d\Pi(Q)}{dQ}=F(Q).$ (51)
The boundary conditions (39) tell that the fictitious particle starts its motion at $\tau=-\infty$ at the potential maximum of $\Pi(Q)$ at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$ (see, however, Section VIII.1). We may present these maximum/minimum conditions as
@@
and opposite inequality for $V=V_{1}$. Note that we have reproduced (46) in the particular case of half-nonlinear transmission line. It would be convenient further on to rewrite (50) as
@@
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged.
For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximated $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
@@
where $\beta$ and $C_{s}$ are constants. We also took $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. In this cases, (49) becomes
@@
$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}},$ (56a)
$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}}.$ (56b)
Note that
$\Pi(Q_{1})-\Pi(Q_{2})=\frac{\beta}{6}\left(Q_{1}-Q_{2}\right)\left(Q_{1}^{2}-Q_{1}Q_{2}+Q_{2}^{2}\right),$
hence
$\beta(Q_{1}-Q_{2})>0.$ (58)
We can also improve the approximation (54) by modifying it to
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}},$ (59)
where $\gamma$ is also assumed to be constant. This modification can be important when $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. In this cases, (49) becomes
where
$Q_{3}\equiv\frac{\beta}{\gamma}+Q_{1}+Q_{2}.$ (61)
VII The ODE which doesn’t contain explicitly the independent variable
We would be looking for the solutions of (55) and (60) in terms of elementary functions. Following the well known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us change gears and instead of (60) consider the generalized damped Helmholtz-Duffing equation kogan4
$x_{\tau\tau}+kx_{\tau}=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right),$ (62)
where $n$, $k$, $\gamma$, $x_{1}$, $x_{3}$ are constants, with the boundary conditions
$\lim_{\tau\to-\infty}x(\tau)=x_{1}^{1/n},\hskip 28.45274pt\lim_{\tau\to+\infty}x(\tau)=0.$ (63)
Equation (62) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $x$ as the new independent variable and
$p=\frac{dx}{d\tau}$ (64)
as the new dependent variable. In the new variables (62) takes the form of Abel equation of the second kind polyanin .
$pp_{x}+kp=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (65)
The boundary conditions in the new variables are
$p\left(x_{1}^{1/n}\right)=p(0)=0.$ (66)
One can easily check up that for $\gamma$ and $k$ connected by the formula
$k^{2}=\frac{\gamma\left[x_{1}+(n+1)x_{3}\right]^{2}}{n+1},$ (67)
the solution of (65) satisfying the boundary conditions (66) is
$p=mx\left(x^{n}-x_{1}\right),\hskip 28.45274ptm=\sqrt{\frac{\gamma}{n+1}}.$ (68)
Substituting $p(x)$ into (64) and integrating we obtain the solution of (62) as
$x(\tau)=\frac{x_{1}^{1/n}.}{\left[\exp\left(nmx_{1}\tau\right)+1\right]^{1/n}}.$ (69)
Now consider the equation
$x_{\tau\tau}+kx_{\tau}=\beta x\left(x^{n}-x_{1}\right)$ (70)
with the boundary conditions (63). We can present (70) as
$x_{\tau\tau}+kx_{\tau}=\beta x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right).$ (71)
Hence we realize that for $k$ and $\beta$ connected by the formula
$k=(n+4)\sqrt{\frac{\beta x_{1}}{2(n+2)}},$ (72)
the solution of (70) satisfying the boundary conditions (66) is
$x(\tau)=\frac{x_{1}^{1/n}.}{\left\{\exp\left[n\sqrt{\frac{\beta x_{1}}{2(n+2)}}\tau\right]+1\right\}^{2/n}}.$ (73)
To additionally illustrate the method of integration used in this Section, in the Appendix B we recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
Let us return to (62) and modify it to
$x_{\tau\tau}+k(1+2\beta x^{n})x_{\tau}=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (74)
Thus we take into account possible nonlinearity of friction. Thus instead of (65) we obtain
$pp_{x}+k(1+2\beta x^{n})p=\gamma x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (75)
Acting as above we obtain that for $k$ and $\gamma$ connected by the formula
$k^{2}=\frac{\gamma[x_{1}+(n+1)x_{3}]^{2}}{\left(1-2\beta x_{3}\right)\left(n+1+2\beta x_{1}\right)},$ (76)
the solution of (75) satisfying the boundary conditions (66) is (68) (same as it was for $\beta=0$), only this time
$m=\sqrt{\frac{\gamma\left(1-2\beta x_{3}\right)}{n+1+2\beta x_{1}}}.$ (77)
Hence the solution of (74) with $\beta\neq 0$ is of the same form as for $\beta=0$ (Eq. (69)), only with the modified $m$.
VIII Back to the transmission line
Now let us return to the particular cases of Eq. (49) (with the boundary conditions (39a)) considered in Section VI. For (60) using the results of the previous Section we claim that when the boundary conditions satisfy the relation
the solution of the equation can be expressed in terms of elementary functions:
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (79)
where
$\psi=\sqrt{\frac{\gamma}{2}}\cdot\frac{Q_{1}-Q_{2}}{T}.$ (80)
For (55) we claim that when the boundary conditions satisfy the relation
$\frac{\tau_{R}}{T}=5\sqrt{\frac{\beta\left(Q_{1}-Q_{2}\right)}{6}},$ (81)
the solution of the equation can be expressed in terms of elementary functions:
@@ VIII Back to the transmission line
$\chi=\sqrt{\frac{\beta(Q_{1}-Q_{2})}{6}}\cdot\frac{1}{T}.$ (83)
Note that neither (78) goes to (81) nor (79) goes to (82) when $\gamma\to 0$. Small change in the equation leads to big change of the condition for the solution to be expressed in elementary functions.
We used previously the linear approximation for $V(Q)$ in the last term in the l.h.s. of (49). Strictly speaking, since we are considering nonlinear $V(Q)$ in the r.h.s. of (49) more consistent would be to treat the same way the l.h.s. Equations. (76), (77) allow us to go one step in this direction, that is to consider $V(Q)$ in the last term in the l.h.s. of (49) in quadratic approximation (for cubic nonlinearity of $V(Q)$). The modification of Eqs. (78), (79) can be easily obtained from Eqs. (76), (77) for $n=1$.
VIII.1 The shocks vs. the kinks
The main subject of the present paper is the shock waves. However, Eq. (49) can describe both the shocks and the kinks kogan4 . To understand the difference between two types of waves note that, strictly speaking, the boundary conditions (39) tell us that the fictitious particle ends its motion at $\tau=+\infty$ at the stationary point $Q_{2}$, which can be either potential minimum or potential maximum. For former case corresponds to the shocks, the latter - to the kinks kogan4 . Equation (55) can describe only the shocks. Equation (60), when the point $-Q_{3}$ is between $Q_{2}$ and $Q_{1}$, describes the kink, when the point is outside of that interval - the shocks.
Equations (78) and (79) are valid both for the kinks and for the shocks, but in these two cases the condition (78) has different meanings. For the kinks it’s the condition of their existence. For the shocks it’s the condition for them to be expressed in elementary functions (actually the condition that the fictitious particle doesn’t oscillate in the vicinity of the stable equilibrium). But, of course, the shocks exist for the general boundary conditions. The same can be said about the shocks described by Eq. (55).
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen, V. Verbuk and A. Volkov for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems and Workshop on Localization and Ergodicity, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A Shock profile for the fully nonlinear transmission line
Let us return to (38). The equation is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (38) either $R_{C}=0$ or $R_{L}=\infty$. In the first case we equate the r.h.s. of (38b) to zero and obtain
$V=U\Phi+\text{const}_{2}.$ (84)
$I=UQ+\text{const}_{1}.$ (85)
If both $R_{C}>0$ and $R_{L}<\infty$, from (38) we may obtain the equation
$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U\Phi-V+\text{const}_{2}}{UQ-I+\text{const}_{1}}.$ (86)
If we manage to solve (86), (3) we can substitute the solution into one of the equations (38) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (55) and (60) describing weak shocks in half-nonlinear transmission line (i.e. Eqs. (78) and (81)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (86) we postpone until later.
Appendix B Solitary wave solution of a generalized KdV equation
Consider a generalized KdV equation drazin